find the slope if it exist
x=-5

Found 2 solutions by Tobiasz, richwmiller:
Answer by Tobiasz(54)

just use y = mx + b

Answer by richwmiller(17219)

You can't apply y=mx+b since y doesn't appear.
since x never changes the slope is undefined.
m=(y-y)/(x-x)